What are the slope and y intercept of the linear equation below?
y=2/3x-1

Answer by stanbon(75887)

What are the slope and y intercept
of the linear equation below?
---------------------------------
y = (2/3)x-1
----------------------
slope = 2/3 ; intercept = -1
---------------------------------
Cheers,
Stan H.
---------------